# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v2

@@ 1.3 Discussion
Previous uniform samplers like $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, it does nothing), to ensure convergence to the target stationary distribution. However, our algorithm does not need this, as our sampler is shown to directly contract towards the target.
Unified framework.
@@
We remark that Theorem 5 places the previously known mixing guarantees for $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework. Existing tight guarantees for $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [LV17]. The known convergence guarantees of $\mathsf{Ball\ walk}$ (see §B for details), namely the mixing time of $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm [$\mathsf{Speedy\ walk}+$rejection sampling]. Here $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step). On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ actually converges to $\pi^{\mathcal{K}}$ without any adjustments and achieves stronger Rényi divergence bounds in the same asymptotic complexity. Our analysis shows that the mixing guarantee is determined by isoperimetric constants of the target (Poincaré or log-Sobolev).
Effective step size.
@@
The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see §B) to keep the rejection probability bounded by a constant. This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$, since the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$ and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{\mathcal{O}}(1/d^{2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
What has really changed?
$\mathsf{In\text{-}and\text{-}Out}$ has clear similarities to both $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$. What then are the changes that allow us to use continuous-time interpolation? One step of $\mathsf{Ball\ walk}$ is [random step ($y\in B_{\delta}(x)$) $+$ Metropolis-filter (accept if $y\in\mathcal{K}$)]. This filtering is an abrupt discrete step, and it is unclear how to control contraction. It could be replaced by a step of $\mathsf{Speedy\ walk}$ ($x\sim\text{Unif}(B_{\delta}(y)\cap\mathcal{K})$). Then, each iteration of $\mathsf{In\text{-}and\text{-}Out}$ can be viewed as a Gaussian version of a $\mathsf{Ball\ walk^{\prime}s\ proposal}+\mathsf{Speedy\ walk}$ algorithm.
@@
How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution. As clarified in Remark 3, one step of (a Gaussian version of) $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow. Therefore, if one can control the isoperimetric constants of the biased distribution along the trajectory of the backward flow, then contraction of $\mathsf{Speedy\ walk}$ toward the biased distribution will follow from the simultaneous backward analysis.
@@ 2 Preliminaries
[…]
To quantify the convergence rate, we introduce some common divergences between distributions.
@@
We state two important lemmas which are needed for our proofs. The first is the data-processing inequality for Rényi divergence and $f$-divergence, given below.
@@ 3 Analysis
We begin this section by proving some introductory lemmas, which will streamline our later exposition.
We first record two fundamental lemmas, which introduce the mathematical formalism for our analysis. The first is the existence of forward and backward heat flows (Lemma 22), which will interpolate each line in Algorithm 1. These flow equations describe how the laws of $Z_{t}$ and $Z_{t}^{\leftarrow}$ in ($\mathsf{FH}$) and ($\mathsf{BH}$) evolve respectively over time.
@@
The forward heat flow equation with initial distribution $\mu_{0}$ is given by
$\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}\,,$
and its backward heat flow equation is given by
$\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\pi^{X}P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{with }\mu_{0}^{\leftarrow}=\mu_{h}\,.$
These admit (weak) solutions on $[0,h]$ for any initial distribution $\mu_{0}$ with $\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\pi^{X}}\leq M<\infty$.
One successful iteration of $\mathsf{In\text{-}and\text{-}Out}$ is exactly the same as the composition of running the forward heat flow and then backward heat flow, both for time $h$.
Lemma 12.
Let $\mu_{k}^{X}$ be the law of the $k$-th iterate $x_{k}$ of $\mathsf{In\text{-}and\text{-}Out}$. If ($\mathsf{FH}$) is initialized with $\mathsf{law}(Z_{0})=\mu_{k}^{X}$, then $\mathsf{law}(Z_{h})=\mu_{k+1}^{Y}$. If ($\mathsf{BH}$) is initialized with $\mathsf{law}(Z_{0}^{\leftarrow})=\mu_{k+1}^{Y}$, then $\mathsf{law}(Z_{h}^{\leftarrow})=\mu_{k+1}^{X}$.
@@
We explicitly show that the forward and backward heat flows indeed interpolate the two discrete steps given in Algorithm 1. For the forward part, we have $Z_{h}=Z_{0}+\zeta$ for $\zeta\sim\mathcal{N}(0,hI_{d})$, so
$\mathsf{law}(Z_{h})=\mathsf{law}(Z_{0})*\mathcal{N}(0,hI_{d})=\mu_{k}^{X}*\mathcal{N}(0,hI_{d})=\mu_{k+1}^{Y}\,.$
Regarding the backward part, it is known from [CCSW22, Lemma 14] that the construction of the time-reversal SDE ensures that $(Z_{h}^{\leftarrow},Z_{0}^{\leftarrow})$ and $(Z_{0},Z_{h})$ have the same joint distribution, when $Z_{0}\sim\pi^{X}$ (and so $Z_{h}\sim\pi^{Y}$). Hence, $\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)=\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$, where the last equality follows from $(Z_{0},Z_{h})\sim\pi$. Since we initialize ($\mathsf{BH}$) with $Z_{0}^{\leftarrow}=y\sim\mu_{k+1}^{Y}$, we have
$\mathsf{law}(Z_{h}^{\leftarrow})=\int\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)\,\mu_{k+1}^{Y}(\mathrm{d}y)=\int\pi^{X|Y}(\cdot|y)\,\mu_{k+1}^{Y}(\mathrm{d}y)=\mu_{k+1}^{X}\,,$
where the last follows from the definition of Line 3. ∎
Our analysis of these objects consists of two parts: (1) demonstrating the mixing of $\mathsf{In\text{-}and\text{-}Out}$, i.e., how many outer iterations are needed to be sufficiently close to the uniform distribution, and (2) quantifying the failure probability and wasted steps in Line 3.
For (1), we collect in §3.1 some important implications of functional inequalities, e.g. the Poincaré and log-Sobolev inequalities, for the uniform distribution. Then in §3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees. To this end, we revisit the proofs for the contraction results of forward and backward heat flows, and then extend them to our constrained setting:
Theorem 13 (Adapted for uniform distributions, [CCSW22]).
Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$. Let $C_{\mathsf{LSI}}$ be the (8) constant of the uniform distribution $\pi^{X}$ over $\mathcal{K}$. Then, for any $q\geq 1$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{2k/q}}\,.$
For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$ and any $q\geq 2$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{2k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\,,\\ (1+h/C_{\mathsf{PI}})^{-2(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\rceil\,.\end{cases}$
@@
Convexity of $\mathcal{K}$ is crucial this time unlike (1). We show in §3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$, and that the expected number of trials per iteration is of order $\log N$, not $N$.
@@
Let $\mathcal{K}$ be any convex body in $\mathbb{R}^{d}$ presented by a well-defined membership oracle, $\pi^{X}$ the uniform distribution over $\mathcal{K}$, and $\mu$ an $M$-warm initial distribution with respect to $\pi^{X}$. For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=\frac{\log\log Z}{2d^{2}\log Z}$ and $N=Z\log^{4}Z=\widetilde{\mathcal{O}}(\nicefrac{{mM}}{{\eta}})$. Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$, and the expected membership queries per iteration is $\mathcal{O}\bigl{(}M\log^{4}\nicefrac{{mM}}{{\eta}}\bigr{)}$.

### v3

@@ 1.3 Discussion
No need to be lazy.  Previous uniform samplers like the $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, it does nothing), to ensure convergence to the target stationary distribution. However, our algorithm does not need this, as our sampler is shown to directly contract towards the target.
@@
Unified framework.  We remark that Theorem 5 places the previously known mixing guarantees for the $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework. Existing tight guarantees for the $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [LV24]. The known convergence guarantees of the $\mathsf{Ball\ walk}$ (see §A for details), namely the mixing time of $\widetilde{O}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm $\mathsf{Speedy\ walk}+$rejection sampling]. Here, the $\mathsf{Speedy\ walk}$ records only the accepted steps of the $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step). On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ actually converges to $\pi^{\mathcal{K}}$ without any adjustments and achieves stronger Rényi divergence bounds in the same asymptotic complexity. Our analysis shows that the mixing guarantee is determined by isoperimetric constants of the target (Poincaré or log-Sobolev).
@@
Effective step size.   The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see §A) to keep the rejection probability bounded by a constant. This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$, since the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$ and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{O}(d^{-2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
[…]
@@
How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with the $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution. One step of (a Gaussian version of) the $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow. Therefore, if one can control the isoperimetric constants of the biased distribution along the trajectory of the backward flow, then contraction of the $\mathsf{Speedy\ walk}$ toward the biased distribution will follow from the simultaneous backward analysis.
Subsequent work.  The ideas and methods of this paper have led to subsequent progress. First, [KZ25] shows that the output guarantee can be strengthened to $\mathcal{R}_{\infty}$ under LSI. Second, the diffusion-based approach is extended to the general logconcave sampling in [KV25b], beyond the uniform distribution. Lastly, the $\mathcal{R}_{\infty}$-warmness requirement has been relaxed to significantly weaker $\mathcal{R}_{\widetilde{O}(1)}$-warmness in [KV25a].
1.4 Related work
Sampling from constrained logconcave distributions is a fundamental task arising in many fields. Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [DFK91]. Since then, this fundamental problem has been studied for over three decades [LS90, LS93, KLS97, LV06, BEL18, BDMP17]. We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers. Below, $\mathcal{K}$ is convex.
Geometric random walks.  We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [LS93, KLS97] and $\mathsf{Hit\text{-}and\text{-}Run}$ [Smi84, Lov99]. The $\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise. In the literature, the $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of $\mathsf{Speedy\ walk}+$ rejection sampling], where the $\mathsf{Speedy\ walk}$ records only the accepted steps of the $\mathsf{Ball\ walk}$ (see §A for details). The step size $\delta$ should be set to $O(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$. [KLS97] showed that the $\mathsf{Ball\ walk}$ needs $\widetilde{O}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$. [LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lvert X-\mathbb{E}X\rvert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $O(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries. Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to the $\mathsf{Ball\ walk}$.
Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding. This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling. In literature, there exists a randomized algorithm [JLLV21] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=O(1)$), with the $\mathsf{Ball\ walk}$ used as a core subroutine. We refer readers to [KZ24] for a streamlined proof, with $\mathsf{In\text{-}and\text{-}Out}$ used as a main sampler. This algorithm takes up $\widetilde{O}(d^{3.5})$ queries in total, and from a well-rounded position with a good warm start, the $\mathsf{Ball\ walk}$ only needs $\widetilde{O}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
Structured samplers.  The aforementioned samplers based on geometric random walks require only access to the membership oracle of the convex body without any additional structural assumptions. The alternate paradigm of geometry-aware sampling attempts to exploit the structure of convex constraints, with the aim of expediting the convergence of the resultant sampling schemes. One common assumption is to make available a self-concordant barrier function $\phi$ which has regularity on its high-order derivatives and blows up when approaching the boundary $\partial\mathcal{K}$. The Hessian of $\phi$ encodes the local geometry of the constraint, and the samplers often work directly with $\nabla^{2}\phi$.
The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [KN12]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment. [KN12] shows that $\mathsf{Dikin\ walk}$ mixes in $O(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query. This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{O}(md)$ iterations of $\mathsf{Dikin\ walk}$ [KV24].
$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [GC11]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter. In the linear-constraint setting above, it requires $O(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [LV18]. This sampler is further analyzed for practical ODE solvers [KLSV23] and for more sophisticated self-concordant barriers [GKV24].
Similarly, $\mathsf{Mirror\ Langevin}$ [ZPFP20, Jia21, AC21, LTVW22] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$. The algorithm can also be metropolized [SWW24]. The best known rate for this algorithm is $\widetilde{O}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
Diffusion-based samplers.  Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [BGHM95, DT12, Dal17, Che25]. While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution. This is clearly impossible to achieve in the constrained setting, so some techniques are required to circumvent this difficulty. These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by incorporating first-order information from $f$.
The first approach for adapting diffusion-based samplers [BEL18, Leh23] iterates a two-step procedure. First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step hh, [^2] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$. and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$. The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\widetilde{O}(\nicefrac{{d^{2}D^{3}}}{{\varepsilon^{4}}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$. Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [GHZ24], and achieves a rate estimate of $\widetilde{O}(\nicefrac{{d}}{{\varepsilon^{10}}})$.
A second approach, suggested by [BDMP17], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\,\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$. It is called Moreau–Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization. This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly. Their final rate estimate is $\widetilde{O}(\nicefrac{{d^{5}}}{{\varepsilon^{6}}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
Observing the prior work integrating diffusion-based sampling with convex constraints, the dependence on the key parameters $d,\varepsilon$, while polynomial, are many orders worse than the rates for zeroth-order samplers such as $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$. In contrast, our analysis not only recovers but in some sense surpasses the known rates for $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$, while harmonizing well with the continuous-time perspective of diffusions.
Proximal schemes for sampling.  The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [LST21] (in turn inspired by Gaussian cooling [CV18]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition). This scheme was given novel interpretations by [CCSW22], which showed that it interpolates the forward and backward heat flows, in the sense defined by [KP23]. The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [Eld13, Che21].
@@ 2 Preliminaries
Unless otherwise specified, we will use $\lvert\cdot\rvert$ for the $2$-norm on $\mathbb{R}^{d}$. We write $a=O(b)$, $a\lesssim b$ to mean that $a\leq cb$ for some universal constant $c>0$. Similarly, we write $a\gtrsim b,a=\Omega(b)$ for $a\geq cb$, while $a=\Theta(b)$ means $a\lesssim b,b\lesssim a$ simultaneously. We will also use $a=\widetilde{O}(b)$ to denote $a=O(b\operatorname{polylog}b)$. Lastly, we will use measure and density interchangeably when there is no confusion.
To quantify the convergence rate, we recall common notions of divergence between distributions.
@@
We state two important lemmas which are needed for our proofs. The first is the data-processing inequality (DPI) for Rényi divergence and $f$-divergence, given below.
@@ 3 Analysis
We begin this section by proving the stationarity of the target $\pi^{X}$.
@@
$\pi^{X}$ is stationary under iterations of $\mathsf{In\text{-}and\text{-}Out}$.
@@
Note that the transition kernel of the forward and backward step is $\pi^{Y|X=\cdot}$ and $\pi^{X|Y=\cdot}$, respectively. Hence, the forward step brings $\pi^{X}$ to $\pi^{Y}$ as seen in
$\int\pi^{Y|X}(\cdot\,|\,x)\,\pi^{X}(x)\,\mathrm{d}x=\int\pi^{X,Y}(x,\cdot)\,\mathrm{d}x=\pi^{Y}\,.$
Also, the backward step brings $\pi^{Y}$ to $\pi^{X}$:
$\int\pi^{X|Y}(\cdot\,|\,y)\,\pi^{Y}(y)\,\mathrm{d}y=\int\pi^{X,Y}(\cdot,y)\,\mathrm{d}y=\pi^{X}\,.$
Therefore, $\pi^{X}$ is stationary under each iteration of $\mathsf{In\text{-}and\text{-}Out}$. ∎
Our analysis for $\mathsf{In\text{-}and\text{-}Out}$ consists of two parts: (1) demonstrating its mixing, i.e., how many outer iterations are needed to be sufficiently close to the uniform distribution, and (2) quantifying the failure probability and wasted steps in Line 3.
For (1), we collect in §3.1 some important implications of functional inequalities, e.g., the Poincaré and log-Sobolev inequalities, for the uniform distribution. Then in §3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees. To this end, we revisit the proofs for the contraction results of forward heat flows in [CCSW22, KO25].
Theorem 12.
Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$ and step size $h>0$. Let $C_{\mathsf{LSI}}$ be the (8) constant of the uniform distribution $\pi^{X}$ over $\mathcal{K}$. Then, for any $q\geq 1$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{k/q}}\,.$
For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$,
$\chi^{2}(\mu^{X}_{k}\mathbin{\|}\pi^{X})\leq\frac{\chi^{2}(\mu^{X}_{0}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{PI}})^{k}}\,.$
Furthermore, for any $q\geq 2$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl(\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr)\,,\\ (1+h/C_{\mathsf{PI}})^{-(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl(\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr)\rceil\,.\end{cases}$
@@
As for (2), convexity of $\mathcal{K}$ is crucial this time unlike (1). We show in §3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$, and that the expected number of trials per iteration is of order $\log N$, not $N$.
@@
Let $\mathcal{K}$ be any convex body in $\mathbb{R}^{d}$ presented by a well-defined membership oracle, $\pi^{X}$ the uniform distribution over $\mathcal{K}$, and $\mu$ an $M$-warm initial distribution with respect to $\pi^{X}$. For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=d^{-2}\frac{\log\log Z}{2\log Z}$ and $N=Z\log^{4}Z=\widetilde{O}(\nicefrac{{mM}}{{\eta}})$. Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$, and the expected membership queries per iteration is $O(M\log^{4}\nicefrac{{mM}}{{\eta}})$.